OpenStudy (anonymous):

infinite series of n^7/7n^6 converge or diverge using ratio test

OpenStudy (anonymous):

I can set it up, but I do not know how to cancel things out after that.

OpenStudy (anonymous):

(n+1)^7/(7(n+1)^6) * 7n^6/n^7

OpenStudy (cwrw238):

shouldnt it be a division? its a ratio test

OpenStudy (cwrw238):

oh i see - sorry

OpenStudy (anonymous):

no problem, I am not sure where to go from here

OpenStudy (cwrw238):

i think this reduces to n+1 / 7 * 7 / n = n+1 / n

OpenStudy (anonymous):

but the answer is that it converges to o, how?

OpenStudy (cwrw238):

hmm - right - i thought a ratio n+1/n means diverging - i'll check it out

OpenStudy (cwrw238):

i dont understand this - the ratio test says that if the 'n+1' term / 'n' term is > 1 then the series diverges - if < 1 then it converges.

OpenStudy (zarkon):

\[\sum_{n=1}^\infty \frac{n^7}{7n^6}\] diverges.

OpenStudy (anonymous):

the series diverges...